OpenStudy (anonymous):

solve I 3x+1I >or equal to 4

OpenStudy (anonymous):

\[3x+1\geq 4\] or \[3x+1\leq -4\] you have to solve them separately. first one gives \[x\geq 1\] second gives \[x\leq -\frac{5}{3}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw